OpenStudy (anonymous):

i need help plz x/6 = 6/24 find x

OpenStudy (phi):

you could first simplify 6/24 can you do that ?

OpenStudy (phi):

to find x, one way is multiply both sides of the equation by 6

OpenStudy (anonymous):

so would i diving 6 and 24

OpenStudy (anonymous):

divide*

OpenStudy (phi):

to simplify 6/24 divide top and bottom by 6 that makes the problem a little easier. what do you get ?

OpenStudy (anonymous):

1/4

OpenStudy (phi):

you would get \[ \frac{x}{6}= \frac{1}{4} \] now multiply both sides by 6, like this \[ \frac{x}{6}\cdot 6= \frac{1}{4} \cdot 6\]

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

and idk

OpenStudy (phi):

you get 6/4 or 3/2 (simplified) or 1.5 (as a decimal)

OpenStudy (phi):

you should know that \[ \frac{x}{6} \cdot \frac{6}{1} = \frac{6}{6} \cdot \frac{x}{1}=x \] because we can change the order fo x*6 to 6*x and 6 divided by 6 is 1

OpenStudy (phi):

you could have done it this way: \[ \frac{x}{6} = \frac{6}{24} \] multiply both sides by 6: \[ \frac{x}{6} \cdot \frac{6}{1}= \frac{6}{24} \cdot \frac{6}{1}\] the left side simplifies to just x on the right you could write 36/24, but it is easier to divide top and bottom by 6 to get 6/4 and then divide top and bottom by 2 to get 3/2 so x= 3/2

OpenStudy (anonymous):

i thought x was supposed to be a whole number

OpenStudy (hunus):

It doesn't have to be

OpenStudy (anonymous):

in this question

OpenStudy (phi):

**i thought x was supposed to be a whole number** it would be nice, and if given the ratio 3/2 to 6 we would re-write it as 1/4. but here x = 3/2 because they want the x that works for this particular ratio x/6 to match 1/4 in "real problems" x does not have to be integer.

OpenStudy (phi):

for example, if you were adjusting the cups of flour in a recipe...